OpenStudy (anonymous):

Your rich uncle is giving you $100,000 on the condition that you will spend exactly half of the remaining amount each day (or as close to half as possible). Use the equation y = 100,000(0.5)^x to find out how much you will spend on day 8. Round to the nearest cent. A. $40,000.00 B. $390.63 C. $988.76

Parth (parthkohli):

The equation is already given. So you just substitute x with 8.

OpenStudy (anonymous):

x = to the amount of days...so plug in 8 for x...

OpenStudy (unklerhaukus):

\[y = 100,000(0.5)^x,\qquad\qquad x=8\] \[y = 100,000(0.5)^8=100,000\times\frac12\times\frac12\times\frac12\times\frac12\times\frac12\times\frac12\times\frac12\times\frac12\] \[y=\frac{100,000}{2^8}\]